# Centered Moments of Weighted One-Level Densities of $GL(2)$ $L$-Functions

### v3

@@ 5.2. Trivial Twisting Character
$=\ \phi(0)^{n}\ \mathbb{E}\left[\sum_{t=0}^{n}\binom{n}{t}1^{n-t}(-1-X)^{t}\right]$
$=\ \phi(0)^{n}\ \mathbb{E}\left[(1-1-X)^{n}\right]$
@@
$=\ \begin{cases}0&n\text{ odd},\\ \phi(0)^{n}(n-1)!!\left(\frac{\sigma_{\phi}^{2}}{\phi(0)^{2}}\right)^{n/2}=(n-1)!!\ \sigma_{\phi}^{n}&n\text{ even},\end{cases}$ (5.30)
which is precisely the $n$th centered moment of a Gaussian distribution with variance $\sigma_{\phi}^{2}$, completing the proof of Theorem 1.1 in this case too. $\Box$
Appendix A Switching a Sum and a Product
Switching the following sum and product is what allows us to carry out the combinatorial analysis in Section 5. We present this crucial, albeit cumbersome, step.
@@ Appendix A Switching a Sum and a Product
For clarity of the mainly combinatorial aspects of this proof, we adopt the notation in Section 4:
$T(q,n,m,\chi)\ :=\ \widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n}\frac{\chi(q)^{m}\log^{n}q}{q^{(n+m)/2}\log^{n}Q_{k,N}}\ \sigma_{1}\left(\gcd\left(r,q^{m}\right)\right).$ (A.1)
Furthermore, we abbreviate “not necessarily distinct” as “n.n.d.”. We would like to prove that
@@
$=\$$\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell})\\ \text{n.n.d.};\\ q_{j}\nmid N;\\ \exists j_{1}<j_{2}:\ q_{j_{1}}=q_{j_{2}}\end{subarray}}\left(T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{2}},n_{j_{2}},m_{j_{2}},\chi)\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell;\\ j\neq j_{1},j_{2}\end{subarray}}T(q_{j},n_{j},m_{j},\chi)\right)$ (A.2)
@@
$\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell})\\ \text{n.n.d};\\ q_{j}\nmid N;\\ \exists j_{1}>j_{2}:\ q_{j_{1}}=q_{j_{2}}\end{subarray}}\left(T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{2}},n_{j_{2}},m_{j_{2}},\chi)\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell;\\ j\neq j_{1},j_{2}\end{subarray}}T(q_{j},n_{j},m_{j},\chi)\right)$
$\leq\$$\sum_{\begin{subarray}{c}1\leq j_{2}\leq\ell;\\ 1\leq j_{1}<j_{2}\end{subarray}}\sum_{\begin{subarray}{c}(q_{1},\dots,q_{j_{2}-1},\\ q_{j_{2}+1},\dots,q_{\ell})\\ \text{n.n.d.};\\ q_{j}\nmid N\end{subarray}}\left(T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{1}},n_{j_{2}},m_{j_{2}},\chi)\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell,\\ j\neq j_{1},j_{2}\end{subarray}}T(q_{j},n_{j},m_{j},\chi)\right).$ (A.3)
@@
$\sum_{\begin{subarray}{c}1\leq j_{2}\leq\ell;\\ 1\leq j_{1}<j_{2}\end{subarray}}\sum_{\begin{subarray}{c}(q_{1},\dots,q_{j_{2}-1},\\ q_{j_{2}+1},\dots,q_{\ell})\\ \text{n.n.d.};\\ q_{j}\nmid N\end{subarray}}\left(T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{1}},n_{j_{2}},m_{j_{2}})\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell,\\ j\neq j_{1},j_{2}\end{subarray}}T(q_{j},n_{j},m_{j},\chi)\right)$ (A.4)
$=\$$\sum_{\begin{subarray}{c}1\leq j_{2}\leq\ell;\\ 1\leq j_{1}<j_{2}\end{subarray}}\left[\left(\sum_{q_{j_{1}}\nmid N}T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{1}},n_{j_{2}},m_{j_{2}},\chi)\right)\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell\\ j\neq j_{1},j_{2}\end{subarray}}\left(\sum_{q_{j}\nmid N}T(q_{j},n_{j},m_{j},\chi)\right)\right].$ (A.5)
@@
$=$ $\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right).$ (A.6)

### v4

@@ 5.2. Trivial Twisting Character
(5.38) $=\ \begin{cases}0&n\text{ odd},\\ \phi(0)^{n}(n-1)!!\left(\frac{\sigma_{\phi}^{2}}{\phi(0)^{2}}\right)^{n/2}=(n-1)!!\ \sigma_{\phi}^{n}&n\text{ even},\end{cases}$
which is precisely the $n$th centered moment of a Gaussian distribution with variance $\sigma_{\phi}^{2}$, completing the proof of Theorem 1.1 in the case of $\chi=\chi_{0}$ trivial too.
Appendix A Switching Sums and Products
Switching the following sum and product is what allows us to carry out the combinatorial analysis in Section 5. We present this crucial, albeit cumbersome, step here. For clarity of the mainly combinatorial aspects of this step, we adopt the notation:
(A.1) $T(q,n,m,\chi)\ :=\ \widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n}\frac{\chi(q)^{m}\log^{n}q}{q^{(n+m)/2}\log^{n}Q_{k,N}}\ S_{r}\left(q^{m}\right).$
Furthermore, we abbreviate “not necessarily distinct” as “n.n.d.”.
@@ Appendix A Switching Sums and Products
We would like to prove that
@@
(A.2) $=\$$\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell})\\ \text{n.n.d.};\\ q_{j}\nmid N;\\ \exists j_{1}<j_{2}:\ q_{j_{1}}=q_{j_{2}}\end{subarray}}\left(T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{2}},n_{j_{2}},m_{j_{2}},\chi)\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell;\\ j\neq j_{1},j_{2}\end{subarray}}T(q_{j},n_{j},m_{j},\chi)\right)$
@@
$\sum_{\begin{subarray}{c}(q_{1},\dots,q_{\ell})\\ \text{n.n.d};\\ q_{j}\nmid N;\\ \exists j_{1}<j_{2}:\ q_{j_{1}}=q_{j_{2}}\end{subarray}}\left(T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{2}},n_{j_{2}},m_{j_{2}},\chi)\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell;\\ j\neq j_{1},j_{2}\end{subarray}}T(q_{j},n_{j},m_{j},\chi)\right)$
(A.3) $\leq\$$\sum_{\begin{subarray}{c}1\leq j_{2}\leq\ell;\\ 1\leq j_{1}<j_{2}\end{subarray}}\sum_{\begin{subarray}{c}(q_{1},\dots,q_{j_{2}-1},\\ q_{j_{2}+1},\dots,q_{\ell})\\ \text{n.n.d.};\\ q_{j}\nmid N\end{subarray}}\left(T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{1}},n_{j_{2}},m_{j_{2}},\chi)\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell,\\ j\neq j_{1},j_{2}\end{subarray}}T(q_{j},n_{j},m_{j},\chi)\right).$
@@
(A.4) $\sum_{\begin{subarray}{c}1\leq j_{2}\leq\ell;\\ 1\leq j_{1}<j_{2}\end{subarray}}\sum_{\begin{subarray}{c}(q_{1},\dots,q_{j_{2}-1},\\ q_{j_{2}+1},\dots,q_{\ell})\\ \text{n.n.d.};\\ q_{j}\nmid N\end{subarray}}\left(T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{1}},n_{j_{2}},m_{j_{2}})\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell,\\ j\neq j_{1},j_{2}\end{subarray}}T(q_{j},n_{j},m_{j},\chi)\right)$
(A.5) $=\$$\sum_{\begin{subarray}{c}1\leq j_{2}\leq\ell;\\ 1\leq j_{1}<j_{2}\end{subarray}}\left[\left(\sum_{q_{j_{1}}\nmid N}T(q_{j_{1}},n_{j_{1}},m_{j_{1}},\chi)\cdot T(q_{j_{1}},n_{j_{2}},m_{j_{2}},\chi)\right)\cdot\prod_{\begin{subarray}{c}1\leq j\leq\ell\\ j\neq j_{1},j_{2}\end{subarray}}\left(\sum_{q_{j}\nmid N}T(q_{j},n_{j},m_{j},\chi)\right)\right].$
@@
(A.6) $=$ $\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right).$